# Flow Matching at Scale: A Machine Learning Framework for Efficient Large-Size Sampling of Many-Body Systems

### v2

@@ I Introduction
However, MCMC encounters critical slowing down as correlation lengths grow near continuous phase transitions and toward the continuum limit; in lattice QCD, topological modes can freeze, inflating autocorrelations and complicating error estimates [8]. Fermionic models further face the sign problem, whose generic resolution is NP-hard [9]. Progress relies on advanced updates [7], tempering/multilevel schemes [10], and ML-assisted samplers; notably, flow-based generative sampling have shown decorrelated proposals and efficiency gains in lattice field theories [11].
@@
Early successes with models like Normalizing Flows (NFs) demonstrated the potential to eliminate critical slowing down for a single set of system parameters [11, 13]. However, the inefficiency of these initial approaches in exploring parameter space is a significant limitation. Studying critical phenomena requires simulations across various parameters (e.g., temperatures), but training a separate, computationally expensive model for each point is impractical [14]. A key advancement to address this has been the development of conditional models, which learn the system’s behavior across a continuous parameter range from sparse training data, enabling efficient interpolation and extrapolation [14, 15]. These models are often implemented with either architecturally rigid NFs, which can suffer from training instabilities like mode collapse [13, 16], or with Diffusion Models (DM) [17, 18, 19], which, while robust, typically rely on a computationally intensive iterative sampling process.
@@ II.2 Loss Function
The loss function is similar to the action of the noninteracting fluid system with integration in the entire condition configuration $\phi$ and the initial Gaussian noise. For the path in Eq. (1), conditional velocity field $\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)$ here is just $(\mathbf{x}_{1}-\mathbf{x}_{0})$ and see the Appendix B for more details.
@@ II.3 Flow of Random Sampling
using a method such as explicit Euler or a higher‐order Runge–Kutta. Then, we can transport the samples from the distribution $p_{0}$ to $p_{1}$. For more training and sampling details, see the Appendix C.
@@ III.1 Parameter-Wide Sampling via a Single Network
The training data, consisting of spin configurations $\mathbf{x}_{1}$, uses angles mapped to the range $0,2\pi)$. The neural network is trained to directly output continuous angular values without an explicit modular arithmetic operation (i.e., no wrapping into $(-\pi,\pi]$ or $0,2\pi)$). This is a valid approach because all relevant physical observables, such as the energy and magnetization components, are calculated using trigonometric functions (e.g., $\cos(\Theta_{i}-\Theta_{j})$ and $\sin(\Theta_{i})$), which are inherently periodic with a period of $2\pi$. The full implementation of our model and the source code for all simulations are made publicly available to ensure reproducibility (see Appendix A).
@@ V Result II: Reliable Extrapolation of System Sizes
A finite-size scaling analysis tests the model’s ability to learn scale-invariant physics. Our primary objective here is to qualitatively demonstrate this emergent capability, rather than to achieve a high-precision estimate of $T_{BKT}$. We acknowledge that spin stiffness is a more rigorous metric for analyzing the BKT transition [26]. However, as discussed in our error analysis (Sec. VI), the model’s predictions for fluctuation-related quantities are inherently less precise, a characteristic compounded by our deliberate use of a limited, low-precision training dataset (1500 configurations per temperature from MCMC). Consequently, the statistical noise in the generated stiffness data, particularly near the critical point, was too significant to permit a reliable scaling analysis. We transparently document this behavior alongside other observables in Appendix D.
@@ VII Result IV: Computational Efficiency for Large-Scale Studies
The primary computational advantage of our Flow Matching approach stems from its ability to generalize across system sizes, a capability traditional MCMC methods lack. We compare the total time cost required for a typical finite-size scaling analysis to quantify the practical implications of this paradigm shift in workflow. For our FM model, this involves a one-time training cost on a small lattice ($32\times 32$), followed by rapid inference on a series of larger lattices. The MCMC method requires running a separate, full simulation for each lattice size. We use our highly optimized, GPU-accelerated MCMC algorithm (see Appendix G) as the baseline for this comparison. All timing benchmarks were performed on a single workstation; detailed hardware and software specifications are provided in Appendix H. The key metric for this evaluation is the Sampling Cost, defined as the average wall-clock time to generate a single statistically independent sample:
@@ VIII.2 U-Net Structure
Mechanistically, the encoder uses an pixel shuffle (space-to-depth) [30] to reduce spatial resolution without discarding information, mapping, for example, each $2\times 2$ block into four channels (see Fig. 12). This preserves translation equivariance and expands the effective receptive field exponentially with depth [31], while convolutions at coarser resolutions still implement local updates on block variables. The decoder applies the inverse pixel shuffle (depth-to-space) to reconstruct high-resolution outputs (see Fig. 12), and skip connections concatenate encoder features with aligned decoder features. This fusion preserves fine-grained locality while enforcing cross-scale consistency, allowing the network to integrate global context without sacrificing the learning of scale-consistent local rules at each resolution. For more architectural and technical details, see the Appendix E.
@@
To empirically validate that the U-Net’s inductive biases are key to its successful size extrapolation, we conducted a comparative study against a Diffusion Transformer (DiT) model [32], which lacks strong, built-in locality. The U-Net significantly outperformed the DiT, especially in the extrapolation task, despite having substantially fewer parameters. This result supports our claim that architectural alignment with the problem’s local nature is critical. The detailed experimental setup, results, and discussion of this comparison are presented in Appendix F.
@@ IX.3 Limitations of Flow Matching
A significant limitation highlighted by our analysis is the model’s reduced accuracy in predicting second-moment observables, which are directly related to physical fluctuations. This is evident in the increased MSE for susceptibility as system size grows (see Fig. 9) and, more strikingly, in the statistical noise observed in the extrapolated spin stiffness, as detailed in Appendix D. This behavior is likely rooted in the nature of the regression-based training objective (Eq. 10), which excels at learning the mean of the conditional velocity field but may under-represent its complete variance. This characteristic, in fact, constrains the direct application of this framework to physical problems where high-precision information about fluctuations, such as calculating critical exponents from specific heat or susceptibility, is paramount. Improving the model’s ability to capture the variance of the target distribution faithfully is therefore a critical challenge for future research. While potential avenues could include exploring alternative loss functions, a more immediate and practical solution is integrating our method into a powerful hybrid workflow. In this paradigm, the configurations generated by the FM model serve as high-quality, decorrelated initial states for a subsequent, brief refinement phase using traditional MCMC. A short MCMC chain can efficiently correct for any residual biases from the generative process, ensuring the final ensemble rigorously converges to the target Boltzmann distribution. Combining initial MCMC seeding, massive-scale FM sampling for interpolation and extrapolation, and a final MCMC refinement preserves the core computational advantage. The Flow Matching stage still acts as a potent accelerator, drastically reducing the simulation time required to obtain a large set of independent samples suitable for precision tasks like accurately determining critical exponents.
@@ X Conclusion and Outlook
XI Acknowledgment
We thank Chi-Ting Ho, Hsuan-Yu Wu, and Po-Chung Chen for valuable discussions. This work is supported by the National Center for Theoretical Sciences, the Higher Education Sprout Project funded by the Ministry of Science and Technology, and the Ministry of Education in Taiwan. DWW is supported under the grant NSTC 113-2112-M-007-036.
Appendix A Code Availability
@@
Appendix B Details of Loss Function
@@
Appendix C Training and Sampling Strategy
@@
Appendix D More Physics Observables in Size Extrapolation
@@
Appendix E Neural Network Architecture
@@
E.1 Core Building Blocks
@@
E.2 Optimization of Neural Network Architecture
@@
Appendix F Comparison between U-Net and DiT Architectures
@@
Appendix G GPU-accelerated Monte Carlo Method
@@
G.1 Parallel Architecture and Data Layout
@@
G.2 Checkerboard Metropolis-Hastings Update
@@
G.3 Parallel Tempering (Replica-Exchange)
@@
Appendix H Information about Computing Resources.

### v3

@@ I Introduction
However, MCMC encounters critical slowing down as correlation lengths grow near continuous phase transitions and toward the continuum limit; in lattice QCD, topological modes can freeze, inflating autocorrelations and complicating error estimates [8]. Fermionic models further face the sign problem, whose generic resolution is NP-hard [9]. Progress relies on advanced updates [7], tempering/multilevel schemes [10], and ML-assisted samplers; notably, flow-based generative sampling have shown decorrelated proposals and efficiency gains in lattice field theories [11, 12, 13, 14, 15].
@@
Early successes with models like Normalizing Flows (NFs) demonstrated the potential to eliminate critical slowing down for a single set of system parameters [11, 17]. However, the inefficiency of these initial approaches in exploring parameter space is a significant limitation. Studying critical phenomena requires simulations across various parameters (e.g., temperatures), but training a separate, computationally expensive model for each point is impractical [14]. A key advancement to address this has been the development of conditional models, which learn the system’s behavior across a continuous parameter range from sparse training data, enabling efficient interpolation and extrapolation [14, 18]. These models are often implemented with either architecturally rigid NFs, which can suffer from training instabilities like mode collapse [17, 19], or with Diffusion Models (DM) [20, 21, 22, 23, 24], which, while robust, typically rely on a computationally intensive iterative sampling process.
@@ II.2 Loss Function
The loss function is similar to the action of the noninteracting fluid system with integration in the entire condition configuration $\phi$ and the initial Gaussian noise. For the path in Eq. (1), conditional velocity field $\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)$ here is just $(\mathbf{x}_{1}-\mathbf{x}_{0})$ and see the Appendix A for more details.
@@ II.3 Flow of Random Sampling
using a method such as explicit Euler or a higher‐order Runge–Kutta. Then, we can transport the samples from the distribution $p_{0}$ to $p_{1}$. For more training and sampling details, see the Appendix B.
@@ III.1 Parameter-Wide Sampling via a Single Network
The training data, consisting of spin configurations $\mathbf{x}_{1}$, uses angles mapped to the range $0,2\pi)$. The neural network is trained to directly output continuous angular values without an explicit modular arithmetic operation (i.e., no wrapping into $(-\pi,\pi]$ or $0,2\pi)$). This is a valid approach because all relevant physical observables, such as the energy and magnetization components, are calculated using trigonometric functions (e.g., $\cos(\Theta_{i}-\Theta_{j})$ and $\sin(\Theta_{i})$), which are inherently periodic with a period of $2\pi$. The full implementation of our model and the source code for all simulations are made publicly available to ensure reproducibility (see Sec. XI).
@@ V Result II: Reliable Extrapolation of System Sizes
A finite-size scaling analysis tests the model’s ability to learn scale-invariant physics. Our primary objective here is to qualitatively demonstrate this emergent capability, rather than to achieve a high-precision estimate of $T_{BKT}$. We acknowledge that spin stiffness is a more rigorous metric for analyzing the BKT transition [31]. However, as discussed in our error analysis (Sec. VI), the model’s predictions for fluctuation-related quantities are inherently less precise, a characteristic compounded by our deliberate use of a limited, low-precision training dataset (1500 configurations per temperature from MCMC). Consequently, the statistical noise in the generated stiffness data, particularly near the critical point, was too significant to permit a reliable scaling analysis. We transparently document this behavior alongside other observables in Appendix C.
@@ VII Result IV: Computational Efficiency for Large-Scale Studies
The primary computational advantage of our Flow Matching approach stems from its ability to generalize across system sizes, a capability traditional MCMC methods lack. We compare the total time cost required for a typical finite-size scaling analysis to quantify the practical implications of this paradigm shift in workflow. For our FM model, this involves a one-time training cost on a small lattice ($32\times 32$), followed by rapid inference on a series of larger lattices. The MCMC method requires running a separate, full simulation for each lattice size. We use our highly optimized, GPU-accelerated MCMC algorithm (see Appendix F) as the baseline for this comparison. All timing benchmarks were performed on a single workstation; detailed hardware and software specifications are provided in Appendix G. The key metric for this evaluation is the Sampling Cost, defined as the average wall-clock time to generate a single statistically independent sample:
@@ VIII.2 U-Net Structure
Mechanistically, the encoder uses an pixel shuffle (space-to-depth) [35] to reduce spatial resolution without discarding information, mapping, for example, each $2\times 2$ block into four channels (see Fig. 12). This preserves translation equivariance and expands the effective receptive field exponentially with depth [36], while convolutions at coarser resolutions still implement local updates on block variables. The decoder applies the inverse pixel shuffle (depth-to-space) to reconstruct high-resolution outputs (see Fig. 12), and skip connections concatenate encoder features with aligned decoder features. This fusion preserves fine-grained locality while enforcing cross-scale consistency, allowing the network to integrate global context without sacrificing the learning of scale-consistent local rules at each resolution. For more architectural and technical details, see the Appendix D.
@@
To empirically validate that the U-Net’s inductive biases are key to its successful size extrapolation, we conducted a comparative study against a Diffusion Transformer (DiT) model [37], which lacks strong, built-in locality. The U-Net significantly outperformed the DiT, especially in the extrapolation task, despite having substantially fewer parameters. This result supports our claim that architectural alignment with the problem’s local nature is critical. The detailed experimental setup, results, and discussion of this comparison are presented in Appendix E.
@@ IX.3 Limitations of Flow Matching
A significant limitation highlighted by our analysis is the model’s reduced accuracy in predicting second-moment observables, which are directly related to physical fluctuations. This is evident in the increased MSE for susceptibility as system size grows (see Fig. 9) and, more strikingly, in the statistical noise observed in the extrapolated spin stiffness, as detailed in Appendix C. This behavior is likely rooted in the nature of the regression-based training objective (Eq. 10), which excels at learning the mean of the conditional velocity field but may under-represent its complete variance. This characteristic, in fact, constrains the direct application of this framework to physical problems where high-precision information about fluctuations, such as calculating critical exponents from specific heat or susceptibility, is paramount. Improving the model’s ability to capture the variance of the target distribution faithfully is therefore a critical challenge for future research. While potential avenues could include exploring alternative loss functions, a more immediate and practical solution is integrating our method into a powerful hybrid workflow. In this paradigm, the configurations generated by the FM model serve as high-quality, decorrelated initial states for a subsequent, brief refinement phase using traditional MCMC. A short MCMC chain can efficiently correct for any residual biases from the generative process, ensuring the final ensemble rigorously converges to the target Boltzmann distribution. Combining initial MCMC seeding, massive-scale FM sampling for interpolation and extrapolation, and a final MCMC refinement preserves the core computational advantage. The Flow Matching stage still acts as a potent accelerator, drastically reducing the simulation time required to obtain a large set of independent samples suitable for precision tasks like accurately determining critical exponents.
@@ X Conclusion and Outlook
XI Code Availability
@@
XII Acknowledgment
We thank Chi-Ting Ho, Hsuan-Yu Wu, Po-Chung Chen, and Lingxiao Wang for valuable discussions and suggestions. This work is supported by the National Center for Theoretical Sciences, the Higher Education Sprout Project funded by the Ministry of Science and Technology, and the Ministry of Education in Taiwan. DWW is supported under the grant NSTC 113-2112-M-007-036.
Appendix A Details of Loss Function
@@
Appendix B Training and Sampling Strategy
@@
Appendix C More Physics Observables in Size Extrapolation
@@
Appendix D Neural Network Architecture
@@
D.1 Core Building Blocks
@@
D.2 Optimization of Neural Network Architecture
@@
Appendix E Comparison between U-Net and DiT Architectures
@@
Appendix F GPU-accelerated Monte Carlo Method
@@
F.1 Parallel Architecture and Data Layout
@@
F.2 Checkerboard Metropolis-Hastings Update
@@
F.3 Parallel Tempering (Replica-Exchange)
@@
Appendix G Information about Computing Resources.